# MMRole: A Comprehensive Framework for Developing and Evaluating Multimodal Role-Playing Agents

### v1

@@ Abstract
Recently, Role-Playing Agents (RPAs) have garnered increasing attention for their potential to deliver emotional value and facilitate sociological research. However, existing studies are primarily confined to the textual modality, unable to simulate humans’ multimodal perceptual capabilities. To bridge this gap, we introduce the concept of Multimodal Role-Playing Agents (MRPAs), and propose a comprehensive framework, MMRole, for their development and evaluation, which comprises a personalized multimodal dataset and a robust evaluation method. Specifically, we construct a large-scale, high-quality dataset, MMRole-Data, consisting of 85 characters, 11K images, and 14K single or multi-turn dialogues. Additionally, we present a robust evaluation method, MMRole-Eval, encompassing eight metrics across three dimensions, where a reward model is trained to score MRPAs with the constructed ground-truth data for comparison. Moreover, we develop the first specialized MRPA, MMRole-Agent. Extensive evaluation results demonstrate the improved performance of MMRole-Agent and highlight the primary challenges in developing MRPAs, emphasizing the need for enhanced multimodal understanding and role-playing consistency. The data, code, and models will be available at https://github.com/YanqiDai/MMRole.
@@ 1 Introduction
The advancement of large language models (LLMs) (Zhao et al., 2023) has significantly catalyzed the rise of Role-Playing Agents (RPAs) (Chen et al., 2024b), which are engineered to emulate specific characters and engage in dialogues with human users or other characters. Unlike AI productivity assistants, RPAs primarily focus on delivering emotional value (Li et al., 2023a; Wang et al., 2023; Shao et al., 2023) and facilitating sociological research (Zhou et al., 2023b; Wang et al., 2024a; Chen et al., 2024a; Gu et al., 2024), where typical applications include emotional companions, NPCs in video games, digital clones, and social simulations.
The primary characteristic of RPAs is their capability to engage in human-like and immersive interactions. However, existing studies in role-playing are primarily confined to the textual modality, which has considerable limitations. In the real-world context, human perception integrates multiple modalities, especially visual and textual, allowing for a more direct and comprehensive understanding of the environment than text alone can provide. Therefore, enhancing RPAs with multimodal capabilities is a crucial next step for conducting more realistic and engaging interactions.
In this paper, we introduce the concept of Multimodal Role-Playing Agents (MRPAs). MRPAs are designed to emulate specific characters and engage in dialogues centered around images, with either human users or other characters. Moreover, we propose MMRole, a comprehensive framework for developing and evaluating MRPAs. As presented in Figure 1, this framework includes two principal components: a large-scale, high-quality dataset and a robust evaluation method for MRPAs.
@@
Dataset Construction: The dataset for MRPAs comprises character profiles, images, and dialogues centered around images. Specifically, we consider three categories of characters: fictional characters, historical and public figures, and hypothetical real-life characters. The profiles of the first two categories are summarized by GPT-4 (Achiam et al., 2023) based on information from Wikipedia or Baidu Baike, while those of the last category are randomly generated by GPT-4. For each character, we utilize distinct generic images from MS-COCO (Lin et al., 2014), and manually collect and annotate various character-related images. Finally, the dialogues are generated by GPT-4 based on the character profiles and images, occurring either between a character and a human user or between two characters. Both the character profiles and the dialogues undergo rigorous manual quality control to ensure accuracy and reliability. Statistically, the MMRole-Data dataset comprises 85 characters, 11K images, and 14K dialogues, yielding 85K training samples and 294 test samples.
Performance Evaluation: On one hand, MRPAs must not only accurately emulate specific characters but also deeply comprehend both visual and textual information. On the other hand, existing methods for evaluating RPAs directly score outputs without a ground truth (Zhou et al., 2023a; Tu et al., 2024), which can lead to unstable scoring criteria without a baseline for comparison. Therefore, we propose MMRole-Eval, a robust evaluation method to stably and comprehensively assess MRPAs, which comprises eight metrics across three dimensions: fundamental conversational skills, multimodal understanding abilities, and role-playing qualities. For each metric, our specialized reward model initially conduct a brief qualitative assessment of the relative performance between the evaluated MRPA and the constructed ground-truth data, followed by assigning a quantitative score pair. The final score of the MRPA is the ratio of the two scores within the score pair. To develop the reward model, we employ GPT-4 to assess various MRPAs and leverage the evaluation trajectories to train our reward model, which renders MMRole-Eval both open-source and cost-effective.
Briefly, our main contributions are four-fold:
@@
    We present MMRole-Eval, a robust evaluation method to stably and comprehensively assess MRPAs, comprising eight metrics across three dimensions. A specialized reward model is trained to score MRPAs with the constructed ground-truth data for comparison.
@@ 2 Related Work
Role-Playing Agents. Recent advancements in large language models (LLMs) (Zhao et al., 2023), such as supervised fine-tuning (Wei et al., 2021) and in-context learning (Brown et al., 2020), have significantly catalyzed the rise of Role-Playing Agents (RPAs) (Chen et al., 2024b), which are interactive AI systems that can emulate designated personas. Specifically, the personas can be categorized into individual characters (Wang et al., 2023; Shao et al., 2023; Wang et al., 2024a; Gu et al., 2024) and groups of people with particular attributes (Li et al., 2023b; Hong et al., 2023; Xu et al., 2023). In this study, we primarily focus on the former.
@@
Existing RPAs that emulate individual characters are developed through either training or prompting LLMs with high-quality character-specific dialogues. In a pioneering study, Chen et al. (2023) extracted all dialogue sessions from original scripts to develop a Harry Potter-specific RPA. Furthermore, Wang et al. (2023), Zhou et al. (2023a), Shao et al. (2023) and Li et al. (2023a) constructed hundreds of characters and more comprehensive datasets of character dialogues. These efforts aimed to develop RPAs that can deliver emotional value to humans. In contrast, Gu et al. (2024) focused on facilitating sociological research. They proposed a character simulation framework that explores the role of language in shaping collective behavior through interactive debate scenarios. However, these studies are primarily confined to the textual modality. Conversely, our MMRole framework is the first to enhance RPAs with multimodal capabilities.
@@
The evaluation of RPAs is also a crucial and challenging research direction. Various methods have been proposed. Specifically, Shen et al. (2023) and Chen et al. (2024a) assessed RPAs with multiple-choice questions. Tu et al. (2024) trained a reward model for scoring without a ground truth. Wang et al. (2024b) evaluated the personality fidelity of RPAs through interviews, scoring without a ground truth by GPT-4. Ng et al. (2024) engaged the acquaintances of the target individuals to distinguish between humans and RPAs. Nevertheless, the high expense of human annotation and the potential instability of scoring without a ground truth pose significant challenges. To address this, we develop a reward model to score RPAs with a ground-truth baseline for comparison.
@@
Large Multimodal Models. Large Multimodal Models (LMMs) are advanced AI systems typically built upon LLMs, designed to integrate and comprehend multiple data modalities, particularly text and images (Yin et al., 2023). A variety of impressive LMMs have been released, including closed-source models with hundreds of billions of parameters like GPT-4V (Achiam et al., 2023), Gemini (Team et al., 2023), and Claude 3 (Anthropic, 2024), and open-source models with tens of billions or billions of parameters like MiniGPT-4 (Zhu et al., 2023), InstructBLIP (Dai et al., 2023), LLaVA (Liu et al., 2023), QWen-VL (Bai et al., 2023), InternVL (Chen et al., 2024c), and Yi-VL (Young et al., 2024). Additionally, various innovative techniques have been explored to enhance the performance of LMMs, such as visual instruction tuning (Liu et al., 2023), mixture of experts (Lin et al., 2024), and multi-task balancing (Dai et al., 2024). LMMs are utilized in many vertical fields, such as healthcare (Li et al., 2024), document understanding (Ye et al., 2023), and GUI navigation (Hong et al., 2024). To further explore their potential, we apply LMMs to role-playing for the first time.
@@ 3 Multimodal Role-Playing Agents
In scenarios where the dialogue partner is a human user without a specific identity, given an image $I$, the profile $P$ of the designated character $C$, and the dialogue context $D$, the MRPA steps into the shoes of the character $C$, responding to the human user pertaining to the image $I$:
$D=[h_{1},m_{1},h_{2},m_{2},\ldots,m_{n}],$ (1)
@@ 4.2 Character Profile Generation
Character profiles are crucial for the role-playing effectiveness of MRPAs, especially for those characters with which MRPAs are not familiar. To facilitate a thorough understanding of the designated characters, our character profiles encompass five core parts: brief introduction, personality, life story, main interpersonal relationships, and catchphrases, which are exampled in Appendix B.
As discussed in Section 4.1, three categories of characters are considered in MMRole. The majority of these characters are English, with a smaller proportion being Chinese. For fictional characters, as well as historical and public figures, the profiles are summarized by GPT-4. The information is sourced from Wikipedia for English characters and Baidu Baike for Chinese characters. For hypothetical real-life characters, the profiles are randomly generated by GPT-4 via a two-stage process to ensure both universality and diversity. Firstly, GPT-4 generates the meta information for all characters through a single API call, including basic information such as names and genders, and brief descriptions of their personalities and backgrounds. We instruct GPT-4 that “The character information should cover as many different situations as possible to reflect the diversity and complexity of human society”. Secondly, GPT-4 expands the meta information of each character to derive the profile. The two-stage generation process is exemplified in Appendix C. Additionally, all character profiles undergo rigorous manual quality control to ensure accuracy and reliability, and are simplified by GPT-4 to adhere to the context length limits of most LMMs.
@@ 4.4 Dialogue Generation and Filtering
As discussed in Section 4.1, three types of dialogue scenarios are introduced in MMRole. Based on the character profiles and images, GPT-4 generates dialogues corresponding to each scenario type. Interestingly, we observe that using the prompt, “You are a dedicated role-playing assistant…Please step into the shoes of {character} from {series}” yields better results than the simpler prompt, “You are {character} from {series}”. We suggest that the training data supplied by OpenAI optimizes GPT-4 to function more effectively as a helpful assistant, rather than as an immersive, human-like character. The prompts for dialogue generation are detailed in Appendix D.
To ensure accuracy and reliability, we manually filter all dialogues using several strategies. Specifically, we remove failed response data, as well as non-Chinese and non-English data. Additionally, we eliminate content that replies in the tone of an AI assistant, meaningless modal words frequently output by GPT-4, action and scene descriptions, and unnecessary explanatory prefixes and suffixes.
@@ 5 MMRole-Eval: Performance Evaluation
As illustrated in Figure 1(b), we propose MMRole-Eval, a robust evaluation method to stably and comprehensively assess MRPAs. In this section, we introduce eight evaluation metrics across three dimensions and the approach for score quantification.
@@ 5.2 Score Quantification
To quantitatively evaluate the performance of RPAs across various metrics, existing methods employ reward models or human annotators to directly score outputs without a ground truth (Zhou et al., 2023a; Tu et al., 2024). However, it may be unstable due to the variability of scoring criteria without a baseline for comparison. Therefore, we propose to develop a more stable reward model. Inspired by the evaluation methods of Vicuna (Chiang et al., 2023) and LLaVA (Liu et al., 2024b), our reward model first conduct a brief qualitative assessment of the relative performance between the evaluated MRPA and the constructed ground-truth data for each metric, followed by assigning a quantitative score pair. The final score of the MRPA is the ratio of the two scores within the score pair.
To develop the reward model, we initially employ GPT-4 to assess various MRPAs across all test samples. For each evaluated MRPA and corresponding test sample, GPT-4 outputs brief assessments and score pairs for all metrics through a single API call. The prompt for GPT-4 scoring is provided in Appendix F. Subsequently, these evaluation trajectories are converted into training and validation data for our reward model. Each evaluation trajectory is segmented into eight samples, with each sample evaluating a distinct metric. Compared to directly applying GPT-4 as the reward model, this approach renders MMRole-Eval both open-source and cost-effective.
@@ 6.1 Statistics of MMRole-Data
Table 1 presents the statistics of the MMRole-Data dataset. Totally, the dataset comprises 85 characters, 11,032 images, and 14,346 dialogues, yielding 85,456 training samples and 294 test samples. Specifically, we construct 72 characters and collect 10,975 images for training and in-distribution testing. For out-of-distribution testing, we additionally construct 13 characters and collect 57 images that differ from those used in the former set. Dividing the data in this manner is significant for evaluating the performance of MRPAs on previously unseen characters and images, thereby assessing its generalization ability. Besides, all constructed characters are listed in Appendix A.
@@
MRPAs Params Open-Source Specialized
GPT-4 Turbo (Achiam et al., 2023) $>$ 100B
Gemini Pro Vision (Team et al., 2023) $>$ 100B
Claude 3 Opus (Anthropic, 2024) $>$ 100B
QWen-VL-Max (Bai et al., 2023) $>$ 100B
LLaVA-NeXT-34B (Liu et al., 2024a) 34B
Yi-VL-34B (Young et al., 2024) 34B
InternVL-Chat-V1.5 (Chen et al., 2024c) 26B
QWen-VL-Chat (Bai et al., 2023) 9B
LLaVA-NeXT-Mistral-7B (Liu et al., 2024a) 7B
Yi-VL-6B (Young et al., 2024) 6B
@@
6.2 Development of MMRole-Agent
We fine-tune the QWen-VL-Chat model (Bai et al., 2023) using 8$\times$×A100 GPUs on the training set of MMRole-Data to develop our specialized MRPA, MMRole-Agent. The learning rate is set to $1e$$-5$, and the training is conducted over 3 epochs. Besides, to understand the detailed character profiles and dialogue history, the model maximum length is set to 3072. Other experimental setup and code remain the same as Bai et al. (2023)’s defaults.
6.3 Evaluated MRPAs
To the best of our knowledge, no specialized MRPA has been developed prior to this work. Therefore, our experiments evaluate MMRole-Agent and various existing general-dialogue LMMs across different parameter scales. As presented in Table 3, we select four well-known closed-source LMMs with over 100 billion parameters (Achiam et al., 2023; Team et al., 2023; Anthropic, 2024; Bai et al., 2023), and six widely-used open-source LMMs with tens of billions or billions of parameters (Liu et al., 2024a; Bai et al., 2023; Young et al., 2024; Chen et al., 2024c). For the closed-source models, we utilize their official APIs to conduct performance evaluations. To ensure fairness, each MRPA is queried with the same prompt, as detailed in Appendix E.
6.4 Reward Model in MMRole-Eval
To develop the reward model, we initially utilize GPT-4 to evaluate various general-dialogue LMMs discussed in Section 6.3 across 294 test samples. Statistically, these evaluation trajectories are converted into a total of 23,520 samples, where 320 samples are reserved for validation, with the rest utilized for training. The validation set includes 20 questions, where the responses of two models are randomly selected for each question, and each response is evaluated on all 8 metrics. Subsequently, we fine-tune another QWen-VL-Chat model. The experimental setup and code are the same as those used for developing MMRole-Agent, except that the model maximum length is set to 4096, and the training is conducted over 10 epochs.
To validate the effectiveness of the reward model, we initially calculate the mean absolute errors (MAEs) between the scores evaluated by the reward model and those evaluated by GPT-4. As illustrated in Table 4, the overall MAE for Reward Model (GPT-4) is less than $0.1$. Furthermore, we engage four human evaluators to compare responses from two MRPAs for each metric on each question in the validation set. Humans’ choices among ‘better’, ‘equal’, and ‘worse’ correspond to the score gaps of $0.4$, 00, and $-0.4$ between the two responses, respectively. The average results from all evaluators are then used to determine the ground-truth score gaps. Subsequently, we calculate the MAEs between the score gaps provided by GPT-4 and the ground truth, as well as those between the score gaps provided by the reward model and the ground truth. In this context, score gaps are capped at 0.4 and -0.4. As shown in Table 4, the overall MAEs for GPT-4 (humans) and Reward Model (humans) are comparable and considerably low. The results indicate that our reward model effectively learns the evaluation abilities of GPT-4 and aligns closely with human evaluators.
Table 4: The validation mean absolute error (MAE) results for the effectiveness of the reward model. ‘Reward Model (GPT-4)’ denotes the scores evaluated by the reward model compared to those evaluated by GPT-4. ‘GPT-4 (humans)’ and ‘Reward Model (humans)’ signify the score gaps provided by GPT-4 and the reward model compared to the ground-truth score gaps provided by humans.
Evaluators IA Flu Coh ITR RA PC KC TC Overall

### v2

@@ Abstract
Recently, Role-Playing Agents (RPAs) have garnered increasing attention for their potential to deliver emotional value and facilitate sociological research. However, existing studies are primarily confined to the textual modality, unable to simulate humans’ multimodal perceptual capabilities. To bridge this gap, we introduce the concept of Multimodal Role-Playing Agents (MRPAs), and propose a comprehensive framework, MMRole, for their development and evaluation, which comprises a personalized multimodal dataset and a robust evaluation approach. Specifically, we construct a large-scale, high-quality dataset, MMRole-Data, consisting of 85 characters, 11K images, and 14K single or multi-turn dialogues. Additionally, we present a robust evaluation approach, MMRole-Eval, encompassing eight metrics across three dimensions, where a reward model is designed to score MRPAs with the constructed ground-truth data for comparison. Moreover, we develop the first specialized MRPA, MMRole-Agent. Extensive evaluation results demonstrate the improved performance of MMRole-Agent and highlight the primary challenges in developing MRPAs, emphasizing the need for enhanced multimodal understanding and role-playing consistency. The data, code, and models are all available. [^1] https://github.com/YanqiDai/MMRole
@@ 1 Introduction
The advancement of large language models (LLMs) (Zhao et al., 2023) has significantly catalyzed the rise of Role-Playing Agents (RPAs) (Chen et al., 2024b), which are engineered to emulate specific characters and engage in dialogues with human users or other characters. Unlike AI productivity assistants, RPAs primarily focus on delivering emotional value (Li et al., 2023a; Wang et al., 2023; Shao et al., 2023) and facilitating sociological research (Zhou et al., 2023b; Wang et al., 2024c; Chen et al., 2024a; Gu et al., 2024), where typical applications include emotional companions, NPCs in video games, digital clones, and social simulations.
The primary characteristic of RPAs is their capability to engage in human-like and immersive interactions. However, existing studies in role-playing are primarily confined to the textual modality, which has considerable limitations. In the real-world context, human perception integrates multiple modalities, especially visual and textual, allowing for a more direct and comprehensive understanding of the environment than text alone could provide. Therefore, enhancing RPAs with multimodal capabilities is a crucial next step for conducting more realistic and engaging interactions.
In this paper, we introduce the concept of Multimodal Role-Playing Agents (MRPAs). MRPAs are designed to emulate specific characters and engage in dialogues centered around images, with either human users or other characters. Furthermore, we propose MMRole, a comprehensive framework for developing and evaluating MRPAs. As presented in Figure 1, this framework includes two principal components: a large-scale, high-quality dataset and a robust evaluation approach for MRPAs.
@@
Dataset Construction: The dataset for MRPAs comprises character profiles, images, and dialogues centered around images. Specifically, we consider three categories of characters: fictional characters, historical and public figures, and hypothetical real-life characters. The profiles of the first two categories are summarized by GPT-4 (Achiam et al., 2023) based on information from Wikipedia or Baidu Baike, while those of the last category are randomly generated by GPT-4. For each character, we utilize distinct generic images from MS-COCO (Lin et al., 2014), and manually collect and annotate various character-related images. Finally, the dialogues are generated by GPT-4 based on the character profiles and images, occurring either between a character and a human user or between two characters. Both the character profiles and the dialogues are subjected to rigorous manual quality control to ensure accuracy and reliability. Statistically, the MMRole-Data dataset comprises 85 characters, 11K images, and 14K dialogues, yielding 85K training samples and 294 test samples.
Performance Evaluation: On one hand, MRPAs must not only accurately emulate specific characters but also deeply comprehend both visual and textual information. On the other hand, existing methods for evaluating RPAs directly score outputs without a ground truth (Zhou et al., 2023a; Tu et al., 2024), which may lead to unstable scoring criteria without a baseline for comparison. Therefore, we propose MMRole-Eval, a robust evaluation approach to stably and comprehensively assess MRPAs, which comprises eight metrics across three dimensions: fundamental conversational skills, multimodal understanding abilities, and role-playing qualities. For each metric, our specialized reward model initially conducts a brief qualitative assessment of the relative performance between the evaluated MRPA and the constructed ground-truth data, followed by assigning a quantitative score pair. The final score of the MRPA is the ratio of the two scores within the score pair. To develop the reward model, we employ GPT-4 to assess various MRPAs and leverage the evaluation trajectories to train our reward model, which renders MMRole-Eval both open-source and cost-effective.
Briefly, our main contributions are three-fold:
@@
    We introduce MMRole-Eval, a robust evaluation approach to stably and comprehensively assess MRPAs, comprising eight metrics across three dimensions. A specialized reward model is trained to score MRPAs with the constructed ground-truth data for comparison.
@@ 2 Related Work
Role-Playing Agents. Recent advancements in large language models (LLMs) (Zhao et al., 2023), such as supervised fine-tuning (Wei et al., 2021) and in-context learning (Brown et al., 2020), have significantly catalyzed the rise of Role-Playing Agents (RPAs) (Chen et al., 2024b), which are interactive AI systems that can emulate designated personas. Specifically, the personas can be categorized into individual characters (Wang et al., 2023; Shao et al., 2023; Wang et al., 2024c; Gu et al., 2024) and groups of people with particular attributes (Li et al., 2023b; Hong et al., 2023; Xu et al., 2023; Zhang et al., 2024). In this study, we primarily focus on the former.
@@
Existing RPAs that emulate individual characters are developed through either training or prompting LLMs with high-quality character-specific dialogues. In a pioneering study, Chen et al. (2023) extracted all dialogue sessions from original scripts to develop a Harry Potter-specific RPA. Furthermore, Wang et al. (2023), Zhou et al. (2023a), Shao et al. (2023) and Li et al. (2023a) constructed hundreds of characters and more comprehensive datasets of character dialogues. These efforts aimed to develop RPAs for delivering emotional value to humans. In contrast, Gu et al. (2024) focused on facilitating sociological research. However, these studies are primarily confined to the textual modality. Conversely, our MMRole framework is the first to enhance RPAs with multimodal capabilities.
@@
The evaluation of RPAs is also a crucial and challenging research direction. Diverse methods have been proposed. Specifically, Shen et al. (2023) and Chen et al. (2024a) assessed RPAs with multiple-choice questions. Tu et al. (2024) trained a reward model for scoring without a ground truth. Wang et al. (2024d) evaluated the personality fidelity of RPAs through interviews, scoring without a ground truth by GPT-4. Ng et al. (2024) engaged the acquaintances of the target individuals to distinguish between humans and RPAs. Wang et al. (2024a) further evaluated RPAs in text-based virtual worlds. However, the high expense of human annotation and the potential instability of scoring without a ground truth pose significant challenges. To address this, we develop a reward model to score RPAs with a ground-truth baseline for comparison.
@@
Large Multimodal Models. Large Multimodal Models (LMMs) are advanced AI systems typically built upon LLMs, designed to integrate and comprehend multiple data modalities, particularly text and images (Yin et al., 2023). A variety of impressive LMMs have been released, including closed-source models with hundreds of billions of parameters like GPT-4V (Achiam et al., 2023), Gemini (Team et al., 2023), and Claude 3 (Anthropic, 2024), and open-source models with tens of billions or billions of parameters like MiniGPT-4 (Zhu et al., 2023), InstructBLIP (Dai et al., 2023), LLaVA (Liu et al., 2024c; a; b; Li et al., 2024a), QWen-VL (Bai et al., 2023; Wang et al., 2024b), InternVL (Chen et al., 2024c), and Yi-VL (Young et al., 2024). Additionally, various techniques have been explored to enhance the performance of LMMs, such as visual instruction tuning (Liu et al., 2024c), mixture of experts (Lin et al., 2024), and multi-task balancing (Dai et al., 2024). LMMs are widely applied in vertical fields, including healthcare (Li et al., 2024b), document understanding (Ye et al., 2023), and GUI navigation (Hong et al., 2024). To further explore their potential, we apply LMMs to role-playing for the first time.
@@ 3 Multimodal Role-Playing Agents
In scenarios where the dialogue partner is a human user without a specific identity, given an image $I$, the profile $P$ of the designated character $C$, and the dialogue context $D$, the MRPA steps into the shoes of the character $C$, responding to the human user about the image $I$:
$D=[h_{1},m_{1},h_{2},m_{2},\ldots,h_{n}],$ (1)
@@ 4.2 Character Profile Generation
Character profiles are crucial for the role-playing effectiveness of MRPAs, especially for those characters with which MRPAs are not familiar. To facilitate a thorough understanding of the designated characters, our character profiles encompass five core parts: brief introduction, personality, life story, main interpersonal relationships, and catchphrases, which are exampled in Appendix D.
As discussed in Section 4.1, three categories of characters are considered in MMRole. The majority of these characters are English, with a smaller proportion being Chinese. For fictional characters, as well as historical and public figures, the profiles are summarized by GPT-4 where the information is sourced from Wikipedia for English characters and Baidu Baike for Chinese characters. Furthermore, for hypothetical real-life characters, the profiles are generated through a two-stage process by GPT-4 to ensure both universality and diversity. Firstly, GPT-4 generates the meta information for all characters in a single API call, including basic details such as names and genders, and brief descriptions of personalities and backgrounds. We instruct GPT-4 that “The character information should cover as many different situations as possible to reflect the diversity and complexity of human society”. Secondly, GPT-4 expands the meta information of each character to derive the profile. The two-stage generation process is exemplified in Appendix E. Additionally, all character profiles undergo rigorous manual quality control to ensure accuracy and reliability, detailed in Appendix F, and are simplified by GPT-4 to adhere to the context length limits of most LMMs.
@@ 4.4 Dialogue Generation and Filtering
As discussed in Section 4.1, three types of dialogue scenarios are introduced in MMRole. Based on the character profiles and images, GPT-4 generates dialogues corresponding to each scenario type. Interestingly, we observe that using the prompt, “You are a dedicated role-playing assistant…Please step into the shoes of {character} from {series}” yields better results than the simpler prompt, “You are {character} from {series}”. We suggest that the training data supplied by OpenAI optimizes GPT-4 to function more effectively as a helpful assistant, rather than as an immersive, human-like character. The prompts for dialogue generation are detailed in Appendix I. To ensure accuracy and reliability, we manually filter all dialogues using several strategies, detailed in Appendix F.
@@ 5 MMRole-Eval: Performance Evaluation
As illustrated in Figure 1(b), we propose MMRole-Eval, a robust evaluation approach to stably and comprehensively assess MRPAs. In this section, we introduce eight evaluation metrics across three dimensions and the approach for score quantification.
@@ 5.2 Score Quantification
To quantitatively evaluate the performance of RPAs across various metrics, existing methods utilize reward models or human annotators to directly score outputs without a ground truth (Zhou et al., 2023a; Tu et al., 2024). However, it may be unstable due to the variability of scoring criteria without a baseline for comparison. Therefore, we propose to develop a more stable reward model. Inspired by the evaluation methods of Vicuna (Chiang et al., 2023) and LLaVA (Liu et al., 2024c), our reward model first conducts a brief qualitative assessment of the relative performance between the evaluated MRPA and the constructed ground-truth data for each metric, followed by assigning a quantitative score pair. The final score of the MRPA is the ratio of the two scores within the score pair.
To develop the reward model, we initially employ GPT-4 to assess various MRPAs across all test samples. For each evaluated MRPA and corresponding test sample, GPT-4 outputs brief assessments and score pairs for all metrics through a single API call. Subsequently, these evaluation trajectories are converted into training and validation data for our reward model. Each evaluation trajectory is segmented into eight samples, with each sample evaluating a distinct metric. The prompts for GPT-4 scoring and reward model scoring are provided in Appendix H. Compared to directly applying GPT-4 as the reward model, this approach renders MMRole-Eval both open-source and cost-effective.
@@ 6.1 Statistics of MMRole-Data
Table 1 presents the statistics of the MMRole-Data dataset. Totally, the dataset comprises 85 characters, 11,032 images, and 14,346 dialogues, yielding 85,456 training samples and 294 test samples. Specifically, we construct 72 characters and collect 10,975 images for training and in-distribution testing. For out-of-distribution testing, we additionally construct 13 characters and collect 57 images that differ from those used in the former set. Dividing the data in this manner is significant for evaluating the performance of MRPAs on previously unseen characters and images, thereby assessing its generalization capabilities. All constructed characters are listed in Appendix C.
@@
MRPAs Version Params Open-Source Specialized
GPT-4 Turbo (Achiam et al., 2023) 2024-04-09 $>$ 100B
Gemini Pro Vision (Team et al., 2023) 2023-12-13 $>$ 100B
Claude 3 Opus (Anthropic, 2024) 2024-02-29 $>$ 100B
QWen-VL-Max (Bai et al., 2023) 2023-12-01 $>$ 100B
LLaVA-NeXT-34B (Liu et al., 2024b) 2024-01-30 34B
Yi-VL-34B (Young et al., 2024) 2024-01-23 34B
InternVL-Chat-V1.5 (Chen et al., 2024c) 2024-04-18 26B
QWen-VL-Chat (Bai et al., 2023) 2023-08-22 9B
LLaVA-NeXT-Mistral-7B (Liu et al., 2024b) 2024-01-30 7B
Yi-VL-6B (Young et al., 2024) 2024-01-23 6B
@@
Table 4: The validation mean absolute error (MAE) results for the effectiveness of the reward model. ‘QWen-VL-Chat (GPT-4)’ and ‘Reward Model (GPT-4)’ denote the scores evaluated by QWen-VL-Chat and the reward model compared to those evaluated by GPT-4. ‘QWen-VL-Chat (humans)’, ‘GPT-4 (humans)’, and ‘Reward Model (humans)’ signify the score gaps provided by QWen-VL-Chat, GPT-4, and the reward model compared to the ground-truth score gaps provided by humans.
Evaluators (Ground Truth) IA Flu Coh ITR RA PC KC TC Overall
QWen-VL-Chat (GPT-4) 0.3776 0.3718 0.3218 0.3561 0.3528 0.4091 0.3794 0.4558 0.3780
@@
6.2 Development of MMRole-Agent
We fine-tune the QWen-VL-Chat model (Bai et al., 2023) using 8$\times$×A100 GPUs on the training set of MMRole-Data to develop our specialized MRPA, MMRole-Agent. Integrating data from different characters and dialogue scenarios for multi-task training can improve the generalization capabilities of MMRole-Agent. The learning rate is set to $1e$$-5$, and the training is conducted over 3 epochs. To accommodate detailed character profiles and dialogue history, the model maximum length is set to 3072. Other experimental setups and codes remain the same as Bai et al. (2023)’s defaults.
6.3 Evaluated MRPAs
To the best of our knowledge, no specialized MRPA has been developed prior to this work. Therefore, our experiments evaluate MMRole-Agent and various existing general-dialogue LMMs across different parameter scales. As presented in Table 3, we select four well-known closed-source LMMs with over 100 billion parameters (Achiam et al., 2023; Team et al., 2023; Anthropic, 2024; Bai et al., 2023), and six widely-used open-source LMMs with tens of billions or billions of parameters (Liu et al., 2024b; Bai et al., 2023; Young et al., 2024; Chen et al., 2024c). For the closed-source models, we utilize their official APIs to conduct performance evaluations. To ensure fairness, each MRPA is queried with the same prompt, as detailed in Appendix G.
6.4 Development and Validation of the Reward Model
To develop the reward model, we initially utilize GPT-4 to evaluate various general-dialogue LMMs discussed in Section 6.3 across 294 test samples. Statistically, these evaluation trajectories are converted into a total of 23,520 samples, where 320 samples are reserved for validation, with the rest utilized for training. The validation set includes 20 questions, where the responses of two models are randomly selected for each question, and each response is evaluated on all 8 metrics. Subsequently, another QWen-VL-Chat model (Bai et al., 2023) is trained to develop the specialized reward model. The experimental setup and code are the same as those used for developing MMRole-Agent, except that the model maximum length is set to 4096, and the training is conducted over 10 epochs.
The scoring success rate of the base model QWen-VL-Chat is only $33.13\%$, whereas those of the reward model and GPT-4 are both $100\%$. To further validate the effectiveness of the reward model, we initially calculate the mean absolute errors (MAEs) of the scores evaluated by QWen-VL-Chat and the reward model compared to those evaluated by GPT-4. The score is random if the model fails to score. As illustrated in Table 4, the overall MAE of QWen-VL-Chat (GPT-4) is remarkably high, whereas that for Reward Model (GPT-4) is less than $0.1$. Furthermore, we engage four human evaluators to compare responses from two MRPAs on each metric for every question in the validation set. Their choices among ‘better’, ‘equal’, and ‘worse’ correspond to the score gaps of $0.4$, 00, and $-0.4$ between the two responses, respectively. The results from all evaluators are averaged to obtain the ground-truth score gaps. In this manner, it is easier for human evaluators to yield more consistent results among individuals than directly scoring the MRPA’s responses. Subsequently, we calculate the MAEs between the score gaps provided by QWen-VL-Chat, GPT-4, and the reward model compared to those provided by human evaluators. In this context, score gaps are capped at 0.4 and -0.4. As shown in Table 4, the overall MAE for QWen-VL-Chat (humans) is significantly high, whereas those for GPT-4 (humans) and Reward Model (humans) are comparable and considerably low. Moreover, we report the root mean squared error (RMSE) and Pearson correlation coefficient (Pearson) results in Appendix J. These results indicate that our specialized reward model effectively learns the evaluation abilities of GPT-4 and aligns closely with human evaluators, significantly superior to the non-specialized QWen-VL-Chat model.
Besides, we assess the internal consistency of MMRole-Eval by Cronbach’s alpha coefficient (Cronbach, 1951), where the score of an evaluated MRPA for a given test query is computed as the average across all metrics. The resulting value of $0.70$ indicates that MMRole-Eval exhibits a moderate level of internal consistency. The relatively lower value can be attributed to the diversity of characters and dialogue scenarios, as individual queries inherently assess distinct aspects.